# Would You Trust an AI Doctor? Building Reliable Medical Predictions with Kernel Dropout Uncertainty

### v1

@@ 4.2 Quantitative Comparisons
One of the defining features of our model is its ability to naturally model uncertainty while leveraging the reliable feature vectors provided by existing pre-trained language models. Unlike traditional methods like multi-class SVM, Naïve Bayes, and various deep learning approaches Minaee et al. (2021), which struggle to provide confidence estimates directly, our model inherently incorporates uncertainty quantification. Consequently, we have observed this leads to significantly better performance compared to these alternatives. Besides that, our model leverages prior distributions to encode existing knowledge about the problem domain. This is especially valuable when your data is limited and models such as SVM Cortes and Vapnik (1995) will tend to struggle. Besides, modelling confidence estimates in SVM is not directly possible because the goal in maximum-margin learning is to find the hyperplane with the largest margin between classes, essentially creating a clear decision boundary. They do not inherently encode confidence information within their output. Models such as Naïve Bayes will tend to struggle in our setting because our model offers a principled way to incorporate prior knowledge and uncertainty.
For instance, consider the quantitative results presented by Afzal et al. on the SOAP dataset. Within this comparable setting, our model achieves performance on par with the best-performing BiLSTM model, but with the crucial added benefit of uncertainty modelling. Moreover, on the ROND dataset, our model consistently outperforms text classification results obtained using established large language models such as ChatGPT, Google BARD/Gemini, and GPT-4 Liu et al. (2024). This trend persists in the MT dataset, where the publicly available best model on Kaggle [^3] https://www.kaggle.com/code/ritheshsreenivasan/clinical-text-classification achieves an accuracy of 0.64, which falls short of our model’s performance of 0.69.
@@ 4.3 Language Models Comparison
To assess how strong comparative models work alongside our novel method, we selected various models, with three of them being specifically trained on medical data. All models underwent fine-tuning for 100 epochs using the Adam optimizer (epsilon=1e-8, learning rate=2e-5) and implemented early stopping based on validation loss.
Bio Bert: Bio Bert Lee et al. (2020) has undergone pre-training on an extensive collection of biomedical domain corpora, which includes PubMed abstracts and PMC full-text articles. Blue Bert: Blue Bert Peng et al. (2019b) trained on preprocessed texts from PubMed. Clinical Bert: By Wang et al. (2023b) initialized from BERT underwent training on a substantial multicenter dataset, featuring a large corpus containing 1.2 billion words encompassing diverse diseases. Bert-base-uncased: By Devlin et al. (2018) consists of 110 million parameters, 12 heads, 768 dimensions, and 12 layers, each with 12 self-attention heads. Xlnet-base-cased: Yang et al. (2019) is a generalized autoregressive pre-training model trained on the English language. RoBERTa-base: Proposed by Liu et al. (2019), is trained in the English language using the Masked Language Modeling (MLM) technique. AlBERT-base-v1: Lan et al. (2019) introduced Albert, which shares its layers across its Transformer. It has the same number of parameters and layers as the Bert model. DistilBERT-base-uncased: Sanh et al. (2019) introduced a smaller and faster version of Bert. It is trained on the same dataset as the Bert model.
@@ 4.4 Evaluation Methodology
We carried out a comprehensive evaluation through a series of experiments. We assessed the models’ performance in a few-shot setting to determine which model performs best in low-resource scenarios and how much we can trust its predictions, considering the limited availability of publicly accessible data in the medical domain Spasic et al. (2020). We performed experiments in scenarios involving zero-shot, five-shot, fifteen-shot, and an 80-20 data split. We conducted experiments using a 5-fold cross-validation strategy and presented the mean values in our experimental results.
We presented the macro precision, recall, F1, and accuracy as our primary evaluation metrics. Moreover, as our novel approach allows transformer models to generate predictions in a probabilistic manner, we employed the Brier Score to assess the calibration of the model’s predicted probabilities. The Brier Score Brier (1950), which reflects the mean squared difference between predicted probabilities and actual outcomes, varies from 0 to 1, where 0 signifies a perfect match. Since our experiments encompass both binary and multi-class classification problems, we have presented the respective equations as follows:
@@
Pre Rec F1 Acc Brier Pre Rec F1 Acc Brier Pre Rec F1 Acc Brier
Bio-BERT 0.793 0.811 0.795 0.804 0.073 0.661 0.714 0.674 0.692 0.096 0.972 0.875 0.914 0.952 0.067
Blue-BERT 0.790 0.792 0.787 0.804 0.070 0.651 0.693 0.665 0.692 0.089 0.970 0.874 0.913 0.951 0.068
ClinicalBERT 0.771 0.782 0.774 0.784 0.080 0.632 0.667 0.645 0.673 0.096 0.973 0.877 0.916 0.954 0.056
BERT 0.763 0.772 0.756 0.765 0.097 0.618 0.629 0.630 0.654 0.108 0.846 0.846 0.846 0.905 0.075
Xlnet 0.592 0.599 0.590 0.608 0.150 0.726 0.558 0.512 0.645 0.135 0.763 0.811 0.786 0.855 0.133
RoBERTa 0.540 0.500 0.354 0.392 0.178 0.418 0.517 0.458 0.628 0.127 0.769 0.816 0.788 0.857 0.100
AlBERT 0.494 0.465 0.286 0.353 0.185 0.457 0.523 0.470 0.688 0.117 0.717 0.787 0.737 0.810 0.152
DistilBERT 0.524 0.432 0.322 0.333 0.200 0.453 0.523 0.485 0.660 0.123 0.767 0.815 0.787 0.856 0.138
@@
Bio-BERT 0.760 0.753 0.752 0.765 0.087 0.597 0.581 0.560 0.630 0.134 0.649 0.728 0.650 0.714 0.159
Blue-BERT 0.718 0.728 0.717 0.725 0.112 0.575 0.602 0.559 0.613 0.135 0.649 0.701 0.632 0.704 0.185
ClinicalBERT 0.760 0.788 0.759 0.765 0.111 0.565 0.607 0.577 0.628 0.129 0.717 0.787 0.737 0.810 0.143
BERT 0.615 0.626 0.590 0.588 0.131 0.396 0.497 0.378 0.435 0.180 0.769 0.816 0.788 0.857 0.114
Xlnet 0.523 0.471 0.476 0.490 0.151 0.399 0.410 0.404 0.510 0.162 0.638 0.662 0.646 0.762 0.103
RoBERTa 0.621 0.406 0.228 0.314 0.188 0.693 0.261 0.228 0.252 0.185 0.925 0.625 0.659 0.857 0.107
AlBERT 0.646 0.392 0.213 0.294 0.188 0.639 0.257 0.210 0.232 0.187 0.621 0.712 0.630 0.701 0.223
DistilBERT 0.385 0.347 0.190 0.255 0.191 0.305 0.251 0.193 0.224 0.189 0.716 0.786 0.737 0.809 0.145
@@
Bio-BERT 0.618 0.532 0.527 0.529 0.166 0.508 0.507 0.479 0.543 0.158 0.717 0.787 0.737 0.810 0.203
Blue-BERT 0.602 0.453 0.462 0.490 0.172 0.490 0.493 0.426 0.485 0.172 0.778 0.721 0.743 0.857 0.172
ClinicalBERT 0.500 0.515 0.486 0.569 0.156 0.547 0.515 0.447 0.504 0.157 0.769 0.816 0.788 0.857 0.151
BERT 0.544 0.454 0.394 0.431 0.171 0.474 0.277 0.213 0.262 0.189 0.947 0.750 0.806 0.904 0.160
Xlnet 0.588 0.425 0.385 0.431 0.183 0.380 0.302 0.305 0.380 0.175 0.717 0.790 0.749 0.830 0.122
RoBERTa 0.587 0.197 0.120 0.176 0.197 0.472 0.250 0.181 0.167 0.191 0.595 0.500 0.460 0.390 0.359
AlBERT 0.537 0.339 0.256 0.333 0.187 0.405 0.256 0.163 0.185 0.188 0.586 0.640 0.533 0.571 0.247
DistilBERT 0.568 0.341 0.193 0.275 0.194 0.168 0.265 0.179 0.212 0.186 0.495 0.480 0.530 0.570 0.407
@@
Bio-BERT 0.306 0.281 0.147 0.235 0.204 0.367 0.253 0.289 0.468 0.177 0.595 0.500 0.560 0.490 0.397
Blue-BERT 0.268 0.311 0.277 0.333 0.186 0.253 0.219 0.223 0.314 0.185 0.475 0.375 0.325 0.343 0.369
ClinicalBERT 0.594 0.237 0.334 0.353 0.181 0.197 0.287 0.216 0.248 0.196 0.805 0.500 0.428 0.523 0.198
BERT 0.035 0.219 0.060 0.137 0.213 0.171 0.252 0.183 0.153 0.202 0.820 0.530 0.447 0.541 0.191
Xlnet 0.441 0.248 0.171 0.216 0.192 0.162 0.197 0.156 0.160 0.195 0.494 0.493 0.332 0.333 0.263
RoBERTa 0.382 0.278 0.174 0.255 0.190 0.159 0.247 0.162 0.144 0.193 0.425 0.647 0.427 0.429 0.258
AlBERT 0.342 0.237 0.132 0.353 0.181 0.302 0.239 0.151 0.179 0.189 0.725 0.625 0.619 0.623 0.222
DistilBERT 0.377 0.250 0.171 0.331 0.184 0.154 0.243 0.175 0.209 0.188 0.600 0.529 0.410 0.389 0.287
Table 1: Quantitative Results
@@ 5 Result and Analysis
To comprehensively demonstrate the effectiveness of our model, we undertook a series of extensive comparative analyses. First, we investigated how our novel model-agnostic layer adeptly manages data uncertainties and bolsters the reliability of diverse deep learning models when integrated with them. This analysis utilized the Brier Score to quantify the level of trust we can place in the models’ predictions. Second, we conducted an in-depth analysis to pinpoint instances where the model exhibits lower certainty. This involved examining the predicted probabilities and Brier Scores for each class, allowing us to identify areas where potential improvement lies.
@@ 5.1 Results Discussion
(b) Brier score for each class under different training scenarios for Bio-BERT in the SOAP dataset
@@
Figure 1: Quantitative reliability analysis
@@ 5.2 Overall Discussion
5.3 Kernel Functions and Beta Priors
We experimented with different values for the Beta priors, $\alpha$ and $\beta$. Ideally, we would automate the inference of these prior parameters by placing priors over them themselves, achieving a fully Bayesian model through posterior inference on the hyperpriors. However, this approach incurs a significant computational burden. Fortunately, research on Bayesian approaches like topic models has shown that fixing the prior values can lead to comparable results as those obtained through posterior inference on the priors Blei et al. (2003); Griffiths et al. (2003); Jameel (2014).
Experimenting with different values for the symmetric Beta priors, we found that setting them to 0.1 led to a 10% decrease in the F1 score for legal judgment prediction. Further reducing the values to 0.001 yielded only marginally better results. The optimal configuration was achieved with symmetric Beta priors of $\alpha$ = 0.0001 and $\beta$ = 0.0001, which we used in all our experiments.
@@ 6 Conclusion
7 Limitations
While our current model performs well, we acknowledge several key limitations that we are actively exploring to improve. Firstly, the potential benefits of developing asymmetric priors remain enticing. Though they could enhance results, they introduce significant computational complexity. Choosing priors in Bayesian modelling is a nuanced discussion, and for our work, we prioritize computational efficiency and mathematical tractability by employing conjugate priors. This decision aligns with a widely recognized challenge in Bayesian inference: its computational demands, especially for complex models and large datasets. Real-time applications and scenarios with limited resources often necessitate addressing this issue. Our approach mitigates this concern by utilizing a simple model structure with symmetric and conjugate priors.
However, like any model, Bayesian approaches are susceptible to data quality issues like biases and outliers. Rigorous data cleaning and preprocessing are paramount for reliable results. While we have not observed overfitting or underfitting in our specific case, the potential remains – a common challenge for Bayesian models, particularly with complex structures and limited data. Therefore, we are actively investigating regularization techniques to further enhance the robustness of our work. By acknowledging these limitations and actively seeking solutions, we aim to continuously improve the accuracy, efficiency, and reliability of our Bayesian model.

### v2

@@ 4.2 Quantitative Comparisons
One of the defining features of our model is its ability to naturally model uncertainty while leveraging the reliable feature vectors provided by existing pre-trained language models. Unlike traditional methods like multi-class SVM, Naïve Bayes, and various deep learning approaches [36], which struggle to provide confidence estimates directly, our model inherently incorporates uncertainty quantification. Consequently, we have observed this leads to significantly better performance compared to these alternatives. Besides that, our model leverages prior distributions to encode existing knowledge about the problem domain. This is especially valuable when your data is limited and models such as SVM [11] will tend to struggle. Besides, modelling confidence estimates in SVM is not directly possible because the goal in maximum-margin learning is to find the hyperplane with the largest margin between classes, essentially creating a clear decision boundary. They do not inherently encode confidence information within their output. Models such as Naïve Bayes will tend to struggle in our setting because our model offers a principled way to incorporate prior knowledge and uncertainty.
@@ 4.3 Language Models Comparison
Our primary comparison is with the recent study [37], where the authors introduce a similar technique utilizing Monte Carlo dropout to assess the reliability of their model’s predictions. While their approach employs the BERT model for feature extraction, we expanded our experiments by exploring other models to offer a more comprehensive perspective. Notably, three of the models we chose were specifically trained on medical data. Our work differs from that of Miok et al. [37] in two key aspects: 1) we have incorporated priors over the hyperparameters in our proposed approach, and 2) we leverage the kernel function. All models underwent fine-tuning for 100 epochs using the Adam optimizer (epsilon=1e-8, learning rate=2e-5) and implemented early stopping based on validation loss.
Bio Bert: Bio Bert [28] has undergone pre-training on an extensive collection of biomedical domain corpora, which includes PubMed abstracts and PMC full-text articles. Blue Bert: Blue Bert [42] trained on preprocessed texts from PubMed. Clinical Bert: By [54] initialized from BERT underwent training on a substantial multicenter dataset, featuring a large corpus containing 1.2 billion words encompassing diverse diseases. Bert-base-uncased: By [13] consists of 110 million parameters, 12 heads, 768 dimensions, and 12 layers, each with 12 self-attention heads. Xlnet-base-cased: [58] is a generalized autoregressive pre-training model trained on the English language. RoBERTa-base: Proposed by [31], is trained in the English language using the Masked Language Modeling (MLM) technique. AlBERT-base-v1: [27] introduced Albert, which shares its layers across its Transformer. It has the same number of parameters and layers as the Bert model. DistilBERT-base-uncased: [46] introduced a smaller and faster version of Bert. It is trained on the same dataset as the Bert model.
4.4 Parameter Settings
We experimented with different values for the Beta priors, $\alpha$ and $\beta$. Ideally, we would automate the inference of these prior parameters by placing priors over them themselves, achieving a fully Bayesian model through posterior inference on the hyperpriors. However, this approach incurs a significant computational burden. Fortunately, research on Bayesian approaches like topic models has shown that fixing the prior values can lead to comparable results as those obtained through posterior inference on the priors [5, 20].
Experimenting with different values for the symmetric Beta priors, we found that setting them to 0.1 led to a 10% decrease in the F1 score for prediction. Further reducing the values to 0.001 yielded only marginally better results. The optimal configuration was achieved with symmetric Beta priors of $\alpha$ = 0.0001 and $\beta$ = 0.0001, which we used in all our experiments.
@@ 4.5 Evaluation Methodology
We carried out a comprehensive evaluation through a series of experiments. We assessed the models’ performance in a few-shot setting to determine which model performs best in low-resource scenarios and how much we can trust its predictions, considering the limited availability of publicly accessible data in the medical domain [51]. We performed experiments in scenarios involving zero-shot, five-shot, fifteen-shot, and an 80-20 data split. We conducted experiments using a 5-fold cross-validation strategy and presented the mean values in our experimental results.
We presented F1, and accuracy as our primary evaluation metrics. Moreover, as our novel approach allows transformer models to generate predictions in a probabilistic manner, we employed Root Mean Square Error (RMSE) and Brier Score to assess the calibration of the model’s predicted probabilities. RMSE helps gauge the extent of error variability or dispersion, offering insights into the model’s uncertainty. The Brier Score [6], which reflects the mean squared difference between predicted probabilities and actual outcomes, varies from 0 to 1, where 0 signifies a perfect match. Since our experiments encompass both binary and multi-class classification problems, we have presented the respective equations as follows:
@@
F1 Acc Brier RMSE F1 Acc Brier RMSE F1 Acc Brier RMSE
                                                                                 Baseline Model [37]
BERT 0.738 0.752 0.099 0.277 0.598 0.641 0.133 0.305 0.759 0.857 0.112 0.334
BioBERT 0.781 0.787 0.096 0.223 0.648 0.679 0.110 0.298 0.879 0.890 0.090 0.301
BlueBERT 0.781 0.796 0.085 0.259 0.641 0.675 0.107 0.293 0.864 0.889 0.093 0.305
ClinicalBERT 0.751 0.763 0.091 0.276 0.638 0.659 0.115 0.301 0.885 0.904 0.077 0.278
XLNet 0.575 0.593 0.179 0.310 0.506 0.624 0.193 0.348 0.775 0.849 0.136 0.303
RoBERTa 0.348 0.401 0.183 0.326 0.451 0.617 0.186 0.344 0.801 0.859 0.098 0.301
AlBERT 0.261 0.338 0.196 0.342 0.487 0.632 0.181 0.342 0.726 0.803 0.187 0.339
DistilBERT 0.310 0.322 0.214 0.358 0.477 0.613 0.191 0.353 0.742 0.837 0.166 0.327
                                                                                 Proposed Approach
BERT 0.756 0.765 0.097 0.263 0.630 0.654 0.108 0.296 0.846 0.905 0.075 0.243
BioBERT 0.795 0.804 0.073 0.213 0.674 0.692 0.096 0.233 0.914 0.952 0.067 0.238
BlueBERT 0.787 0.804 0.070 0.219 0.665 0.692 0.089 0.238 0.913 0.951 0.068 0.240
ClinicalBERT 0.774 0.784 0.080 0.248 0.645 0.673 0.096 0.299 0.916 0.954 0.056 0.221
XLNet 0.590 0.608 0.150 0.296 0.512 0.645 0.135 0.306 0.786 0.855 0.133 0.291
RoBERTa 0.354 0.392 0.178 0.303 0.458 0.628 0.127 0.315 0.788 0.857 0.100 0.275
AlBERT 0.286 0.353 0.185 0.315 0.470 0.688 0.117 0.319 0.737 0.810 0.152 0.331
DistilBERT 0.322 0.333 0.200 0.320 0.485 0.660 0.123 0.323 0.787 0.856 0.138 0.302
@@
                                                                                 Baseline Model [37]
BERT 0.563 0.553 0.158 0.379 0.383 0.411 0.205 0.399 0.690 0.754 0.179 0.423
BioBERT 0.732 0.742 0.104 0.352 0.532 0.597 0.177 0.381 0.661 0.711 0.160 0.401
BlueBERT 0.697 0.702 0.136 0.361 0.538 0.596 0.181 0.384 0.633 0.701 0.187 0.410
ClinicalBERT 0.726 0.741 0.133 0.356 0.549 0.602 0.163 0.375 0.728 0.806 0.170 0.413
XLNet 0.477 0.493 0.163 0.382 0.415 0.498 0.207 0.403 0.642 0.759 0.116 0.385
RoBERTa 0.210 0.309 0.213 0.415 0.207 0.249 0.233 0.418 0.633 0.843 0.128 0.391
AlBERT 0.203 0.275 0.217 0.426 0.201 0.225 0.231 0.415 0.628 0.703 0.225 0.450
DistilBERT 0.185 0.233 0.237 0.441 0.201 0.215 0.239 0.424 0.729 0.798 0.151 0.405
                                                                                 Proposed Approach
BERT 0.590 0.588 0.131 0.349 0.378 0.435 0.180 0.372 0.788 0.857 0.114 0.419
BioBERT 0.752 0.765 0.087 0.319 0.560 0.630 0.134 0.345 0.650 0.714 0.159 0.398
BlueBERT 0.717 0.725 0.112 0.336 0.559 0.613 0.135 0.349 0.632 0.704 0.185 0.404
ClinicalBERT 0.759 0.765 0.111 0.325 0.577 0.628 0.129 0.325 0.737 0.810 0.143 0.407
XLNet 0.476 0.490 0.151 0.365 0.404 0.510 0.162 0.381 0.646 0.762 0.103 0.368
RoBERTa 0.228 0.314 0.188 0.389 0.228 0.252 0.185 0.388 0.659 0.857 0.107 0.387
AlBERT 0.213 0.294 0.188 0.405 0.210 0.232 0.187 0.385 0.630 0.701 0.223 0.441
DistilBERT 0.190 0.255 0.191 0.416 0.193 0.224 0.189 0.394 0.737 0.809 0.145 0.402
@@
                                                                                 Baseline Model [37]
BERT 0.340 0.352 0.186 0.429 0.205 0.258 0.214 0.435 0.781 0.887 0.163 0.390
BioBERT 0.496 0.499 0.177 0.415 0.441 0.512 0.196 0.411 0.711 0.806 0.212 0.382
BlueBERT 0.438 0.472 0.193 0.430 0.405 0.469 0.203 0.405 0.726 0.836 0.193 0.395
ClinicalBERT 0.459 0.533 0.171 0.406 0.419 0.472 0.194 0.403 0.776 0.846 0.164 0.391
XLNet 0.371 0.426 0.198 0.469 0.274 0.281 0.219 0.441 0.728 0.813 0.149 0.377
RoBERTa 0.141 0.189 0.229 0.497 0.172 0.150 0.237 0.502 0.473 0.410 0.387 0.491
AlBERT 0.219 0.310 0.221 0.488 0.152 0.169 0.231 0.494 0.531 0.568 0.251 0.436
DistilBERT 0.182 0.248 0.227 0.501 0.159 0.197 0.225 0.487 0.518 0.553 0.410 0.501
                                                                                 Proposed Approach
BERT 0.394 0.431 0.171 0.403 0.213 0.262 0.189 0.397 0.806 0.904 0.160 0.377
BioBERT 0.527 0.529 0.166 0.395 0.479 0.543 0.158 0.372 0.737 0.810 0.203 0.382
BlueBERT 0.462 0.490 0.172 0.401 0.426 0.485 0.172 0.375 0.743 0.857 0.172 0.381
ClinicalBERT 0.486 0.569 0.156 0.373 0.447 0.504 0.157 0.361 0.788 0.857 0.151 0.386
XLNet 0.385 0.431 0.183 0.415 0.305 0.380 0.175 0.406 0.749 0.830 0.122 0.354
RoBERTa 0.120 0.176 0.197 0.436 0.181 0.167 0.191 0.437 0.460 0.390 0.359 0.486
AlBERT 0.256 0.333 0.187 0.472 0.163 0.185 0.188 0.429 0.533 0.571 0.247 0.429
DistilBERT 0.193 0.275 0.194 0.497 0.179 0.212 0.186 0.422 0.530 0.570 0.407 0.500
@@
                                                                                 Baseline Model [37]
BERT 0.157 0.148 0.219 0.485 0.152 0.138 0.227 0.447 0.438 0.522 0.298 0.514
BioBERT 0.129 0.211 0.207 0.457 0.251 0.431 0.203 0.438 0.553 0.486 0.422 0.650
BlueBERT 0.263 0.317 0.201 0.463 0.205 0.288 0.219 0.442 0.331 0.355 0.385 0.563
ClinicalBERT 0.316 0.323 0.200 0.452 0.201 0.237 0.221 0.444 0.410 0.499 0.223 0.505
XLNet 0.173 0.219 0.221 0.487 0.158 0.161 0.220 0.443 0.434 0.331 0.312 0.613
RoBERTa 0.185 0.261 0.218 0.486 0.143 0.141 0.217 0.459 0.404 0.416 0.279 0.512
AlBERT 0.131 0.339 0.216 0.503 0.144 0.159 0.215 0.488 0.520 0.539 0.231 0.511
DistilBERT 0.166 0.290 0.221 0.505 0.153 0.182 0.233 0.495 0.398 0.401 0.276 0.509
                                                                                 Proposed Approach
BERT 0.170 0.137 0.213 0.467 0.183 0.153 0.202 0.435 0.447 0.541 0.191 0.503
BioBERT 0.147 0.235 0.204 0.436 0.289 0.468 0.177 0.399 0.560 0.490 0.397 0.622
BlueBERT 0.277 0.333 0.186 0.447 0.223 0.314 0.185 0.413 0.325 0.343 0.369 0.551
ClinicalBERT 0.334 0.353 0.181 0.409 0.216 0.248 0.196 0.426 0.428 0.523 0.198 0.501
XLNet 0.171 0.216 0.192 0.481 0.156 0.160 0.195 0.422 0.332 0.333 0.263 0.556
RoBERTa 0.174 0.255 0.190 0.479 0.162 0.144 0.193 0.418 0.427 0.429 0.258 0.506
AlBERT 0.132 0.353 0.181 0.491 0.151 0.179 0.189 0.416 0.619 0.623 0.222 0.505
DistilBERT 0.171 0.331 0.184 0.501 0.175 0.209 0.188 0.419 0.410 0.389 0.287 0.510
@@ 5 Result and Analysis
To comprehensively demonstrate the effectiveness of our model, we undertook a series of extensive comparative analyses. First, we investigated how our novel model-agnostic layer adeptly manages data uncertainties and bolsters the reliability of diverse deep learning models when integrated with them. This analysis utilized the RMSE and Brier Score to quantify the level of trust we can place in the models’ predictions. Second, we conducted an in-depth analysis to pinpoint instances where the model exhibits lower certainty. This involved examining the predicted probabilities and Brier Scores for each class, allowing us to identify areas where potential improvement lies.
@@ 5.1 Results Discussion
(b) Brier score for each class under different training scenarios for Bio-BERT in the SOAP dataset.
@@
Figure 1: Uncertainty analysis across multiple datasets.